The Better Business Bureau received the most complaints from consumers about cell-phone poviders and cable/satellite TV provides in 2009. These two industries accounted for a total of 70,093 complaints, with cable/satellite TV providers receiving 4861 fewer complaints than cell-phone providers (Source: Better Business Bureau). How many complaints did each industry receive?
Answer by JulietG(1812)

Let C stand for cellphone providers, and T stand for TV providers.
C + T = 70,093
T = C-4861
Take the value of T from the bottom equation and put it into the top equation.
C + T = 70,093
C + (C-4861) = 70,093
Add
2C - 4861 = 70,093
Add 4861 to each side
2C = 74,954
Divide each side by 2
C = 37,477
.
If C = 37,477, then T = 37,477-4861, or 32,616.
Prove it by adding them together.
37,477 + 32,616 = 70,093
Success!